OpenStudy (help_people):

Which shows 602 − 402 being evaluated using the difference of squares method? 602 − 402 = (3600 + 1600)(3600 − 1600) = 10,400,000 602 − 402 = 3600 − 1600 = 2,000 602 − 402 = (60 − 40)2 = 202 = 400 602 − 402 = (60 + 40)(60 − 40) = (100)(20) = 2,000

OpenStudy (help_people):

@amistre64

OpenStudy (anonymous):

do you mean: evaluate \(60^2-40^2\) using the difference of squares method?

OpenStudy (anonymous):

if so, you can factor...\[60^2-40^2=\left( 60-40 \right)\left( 60+40 \right)\]

OpenStudy (help_people):

so d?

OpenStudy (help_people):

@pgpilot326

OpenStudy (anonymous):

got confidence?

OpenStudy (help_people):

yes